OpenStudy (anonymous):

The area of a circle is about 7 cm squared. By how many times will the area increase if the radius of the circle is tripled?

OpenStudy (anonymous):

help please?(:

OpenStudy (anonymous):

7 *3*3

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

explain further pleazz

OpenStudy (anonymous):

i liked your help Hero

hero (hero):

Okay, let me post the complete solution

OpenStudy (anonymous):

Thank you

hero (hero):

We know: \[A = \pi r^2\] We also know: \[7 = \pi r^2\] Now, we need to find the radius: \[r = \sqrt{\frac{7}{\pi}}\]

hero (hero):

Afterwards we need to multiply that by 3, and re-insert it into the area of a circle formula:

OpenStudy (anonymous):

i dont quite understand how we find the radius

hero (hero):

Do you know how to isolate r in the following equation \[7 = \pi r^2 \]?

OpenStudy (anonymous):

you divide by pi right?

hero (hero):

Okay, if you did that you would end up with \[\frac{7}{\pi} = r^2\] We have r^2, but we still need to isolate r

hero (hero):

What can we do to both sides to isolate r?

OpenStudy (anonymous):

so we square root it right?

hero (hero):

Yes, and that is how we end up with \[\sqrt{\frac{7}{\pi}} = r\]

OpenStudy (anonymous):

oh isee..then from there what is our next step?

hero (hero):

You found r, but we have to triple it. Once we triple it, we place it back into the area formula, then find the value of A

hero (hero):

Notice that r can be re-written as: \[\frac{\sqrt{7}}{\sqrt{\pi}}\]

hero (hero):

Now we have to multiply this by 3

OpenStudy (anonymous):

how do you multiply sumthin like that by 3?

hero (hero):

When we do this we get: \[r = \frac{3 \sqrt{7}}{\sqrt{\pi}}\]

hero (hero):

Now we place that back into the area formula: \[A = \pi (\frac{3 \sqrt{7}}{\sqrt{\pi}})^2\]

hero (hero):

Amazingly, we get A = 63 Looks like @matricked shortcut worked

OpenStudy (anonymous):

how dipe that last equation in the calculator?

OpenStudy (anonymous):

not dipe but *type

hero (hero):

Depends on what kind of calc you have. I have TI-Nspire, the only calc I would recommend

OpenStudy (anonymous):

i have some sort od scientic one that can handle this but i dont get how to multiply out the radicals?

OpenStudy (anonymous):

By the way the answer is not 63 cuz its not one of the options

hero (hero):

What are the answer choices? Two people use two different methods and come up with the same answer. I'm pretty sure it is correct.

hero (hero):

I performed every step correctly.

OpenStudy (anonymous):

Reread the original problem question cuz I think yo guys forgot something. The answers are 1.5, 3, 6 and 9

hero (hero):

Oh, I see. 63/7 = 9

hero (hero):

It asked "by how much" does the area increase. It increased from 7 to 63 7 x 9 = 63, therefore it increased by a factor of 9

OpenStudy (anonymous):

thank you i got it

hero (hero):

yw

OpenStudy (anonymous):

can u help me with a few more other problems? youe explain it very good

hero (hero):

You assume I have time to explain every step